# Symmetry and Topology in the Non-Hermitian Kitaev chain

### v2

@@ A.4 Adjoint particle hole symmetry (PHS)
$-H(-k)$ $=\left(\begin{array}[]{cc}t_{1}e^{ik}+t_{2}e^{-ik}+2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu\end{array}\right)$ (A62)
$H(k)^{*}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&-2i\Delta_{2}e^{-i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{i\varphi_{1}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (A65)
@@
$-H^{\dagger}(-k)$ $=\left(\begin{array}[]{cc}t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}&2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ -2i\Delta_{2}e^{-i\varphi_{2}}\sin k&-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}\end{array}\right)$ (A68)
$H^{\dagger}(k)^{*}$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&-2i\Delta_{1}e^{-i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{i\varphi_{2}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu^{*}\end{array}\right)$ (A71)
@@
$\quad+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (A74)
@@
$\quad+4t_{1}^{*}t_{2}^{*}+8\Delta_{1}\Delta_{2}e^{i\varphi_{1}-i\varphi_{2}}\sin^{2}k$ (A75)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (A76)
@@
$\quad+4\Delta_{2}^{2}\sin^{2}k+4\Delta_{1}^{2}\sin^{2}k+8|\mu|^{2}$ (A77)
@@ A.5 Chiral symmetry (CS)
$-H(k)$ $=\left(\begin{array}[]{cc}t_{1}e^{-ik}+t_{2}e^{ik}+2\mu&-2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{1}}\sin k&-t_{1}e^{ik}-t_{2}e^{-ik}-2\mu\end{array}\right)$ (A81)
$H(k)^{\dagger}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ -2i\Delta_{2}e^{-i\varphi_{2}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (A84)
@@
$-H(k)^{\dagger}$ $=\left(\begin{array}[]{cc}t_{1}^{*}e^{ik}+t_{2}^{*}e^{-ik}+2\mu^{*}&-2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{2}}\sin k&-t_{1}^{*}e^{-ik}-t_{2}^{*}e^{ik}-2\mu^{*}\end{array}\right)$ (A87)
$H(k)$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (A90)
@@
$\text{Tr}\big[-H(k)\big]$ $=(t_{1}-t_{2})e^{-ik}+(t_{2}-t_{1})e^{ik}$ (A91)
$\text{Tr}\big[H(k)^{\dagger}\big]$ $=(t_{1}^{*}-t_{2}^{*})e^{-ik}+(t_{2}^{*}-t_{1}^{*})e^{ik}$ (A92)
@@
$\text{Tr}\big[H(k)^{2}\big]$ $=(t_{1}^{2}+t_{2}^{2})e^{i2k}+(t_{1}^{2}+t_{2}^{2})e^{-i2k}+4\mu(t_{1}+t_{2})e^{ik}+4\mu(t_{1}+t_{2})e^{-ik}$ (A93)
$+4\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (A94)
$\text{Tr}\Big(\big[H(k)^{\dagger}\big]^{2}\Big)$ $=(t_{1}^{*2}+t_{2}^{*2})e^{i2k}+(t_{1}^{*2}+t_{2}^{*2})e^{-i2k}+4\mu(t_{1}^{*}+t_{2}^{*})e^{ik}+4\mu(t_{1}^{*}+t_{2}^{*})e^{-ik}$ (A95)
$+4\mu^{2}+4t_{1}^{*}t_{2}^{*}+8\Delta_{1}\Delta_{2}e^{i\varphi_{1}-i\varphi_{2}}\sin^{2}k$ (A96)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (A97)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (A98)
@@
$\varphi_{1}$ $=\varphi_{2}$ (A99)
A.6 Adjoint chiral symmetry (CS)
@@
$-H(k)$ $=\left(\begin{array}[]{cc}t_{1}e^{-ik}+t_{2}e^{ik}+2\mu&-2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{1}}\sin k&-t_{1}e^{ik}-t_{2}e^{-ik}-2\mu\end{array}\right)$ (A102)
$H(k)$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (A105)
@@
$-H(k)^{\dagger}$ $=\left(\begin{array}[]{cc}t_{1}^{*}e^{ik}+t_{2}^{*}e^{-ik}+2\mu^{*}&-2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{2}}\sin k&-t_{1}^{*}e^{-ik}-t_{2}^{*}e^{ik}-2\mu^{*}\end{array}\right)$ (A108)
$H(k)^{\dagger}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ -2i\Delta_{2}e^{-i\varphi_{2}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (A111)
@@
$\text{Tr}\big[H(k)^{2}\big]$ $=\text{Tr}\big[\{-H(k)\}^{2}\big]$ (A114)
$\text{Tr}\big[H(k)H(k)^{\dagger}\big]$ $=\text{Tr}\big[\{-H(k)\}\{-H^{\dagger}(k)\}\big]$ (A115)
@@
$\forall$ $\varphi_{1},\,\varphi_{2},\,\mu$ (A116)

### v3

@@ B.4 Adjoint particle hole symmetry (PHS)
$-H(-k)$ $=\left(\begin{array}[]{cc}t_{1}e^{ik}+t_{2}e^{-ik}+2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu\end{array}\right)$ (B63)
$H(k)^{*}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&-2i\Delta_{2}e^{-i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{i\varphi_{1}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (B66)
@@
$-H^{\dagger}(-k)$ $=\left(\begin{array}[]{cc}t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}&2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ -2i\Delta_{2}e^{-i\varphi_{2}}\sin k&-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}\end{array}\right)$ (B69)
$H^{\dagger}(k)^{*}$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&-2i\Delta_{1}e^{-i\varphi_{1}}\sin k\\ 2i\Delta_{2}e^{i\varphi_{2}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu^{*}\end{array}\right)$ (B72)
@@
$\quad+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (B75)
@@
$\quad+4t_{1}^{*}t_{2}^{*}+8\Delta_{1}\Delta_{2}e^{i\varphi_{1}-i\varphi_{2}}\sin^{2}k$ (B76)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (B77)
@@
$\quad+4\Delta_{2}^{2}\sin^{2}k+4\Delta_{1}^{2}\sin^{2}k+8|\mu|^{2}$ (B78)
@@ B.5 Chiral symmetry (CS)
$-H(k)$ $=\left(\begin{array}[]{cc}t_{1}e^{-ik}+t_{2}e^{ik}+2\mu&-2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{1}}\sin k&-t_{1}e^{ik}-t_{2}e^{-ik}-2\mu\end{array}\right)$ (B82)
$H(k)^{\dagger}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ -2i\Delta_{2}e^{-i\varphi_{2}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (B85)
@@
$-H(k)^{\dagger}$ $=\left(\begin{array}[]{cc}t_{1}^{*}e^{ik}+t_{2}^{*}e^{-ik}+2\mu^{*}&-2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{2}}\sin k&-t_{1}^{*}e^{-ik}-t_{2}^{*}e^{ik}-2\mu^{*}\end{array}\right)$ (B88)
$H(k)$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (B91)
@@
$\text{Tr}\big[-H(k)\big]$ $=(t_{1}-t_{2})e^{-ik}+(t_{2}-t_{1})e^{ik}$ (B92)
$\text{Tr}\big[H(k)^{\dagger}\big]$ $=(t_{1}^{*}-t_{2}^{*})e^{-ik}+(t_{2}^{*}-t_{1}^{*})e^{ik}$ (B93)
@@
$\text{Tr}\big[H(k)^{2}\big]$ $=(t_{1}^{2}+t_{2}^{2})e^{i2k}+(t_{1}^{2}+t_{2}^{2})e^{-i2k}+4\mu(t_{1}+t_{2})e^{ik}+4\mu(t_{1}+t_{2})e^{-ik}$ (B94)
$+4\mu^{2}+4t_{1}t_{2}+8\Delta_{1}\Delta_{2}e^{i\varphi_{2}-i\varphi_{1}}\sin^{2}k$ (B95)
$\text{Tr}\Big(\big[H(k)^{\dagger}\big]^{2}\Big)$ $=(t_{1}^{*2}+t_{2}^{*2})e^{i2k}+(t_{1}^{*2}+t_{2}^{*2})e^{-i2k}+4\mu(t_{1}^{*}+t_{2}^{*})e^{ik}+4\mu(t_{1}^{*}+t_{2}^{*})e^{-ik}$ (B96)
$+4\mu^{2}+4t_{1}^{*}t_{2}^{*}+8\Delta_{1}\Delta_{2}e^{i\varphi_{1}-i\varphi_{2}}\sin^{2}k$ (B97)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (B98)
@@
$\quad+4\Delta_{1}^{2}\sin^{2}k+4\Delta_{2}^{2}\sin^{2}k+8|\mu|^{2}$ (B99)
@@
$\varphi_{1}$ $=\varphi_{2}$ (B100)
B.6 Adjoint chiral symmetry (CS)
@@
$-H(k)$ $=\left(\begin{array}[]{cc}t_{1}e^{-ik}+t_{2}e^{ik}+2\mu&-2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{1}}\sin k&-t_{1}e^{ik}-t_{2}e^{-ik}-2\mu\end{array}\right)$ (B103)
$H(k)$ $=\left(\begin{array}[]{cc}-t_{1}e^{-ik}-t_{2}e^{ik}-2\mu&2i\Delta_{2}e^{i\varphi_{2}}\sin k\\ -2i\Delta_{1}e^{-i\varphi_{1}}\sin k&t_{1}e^{ik}+t_{2}e^{-ik}+2\mu\end{array}\right)$ (B106)
@@
$-H(k)^{\dagger}$ $=\left(\begin{array}[]{cc}t_{1}^{*}e^{ik}+t_{2}^{*}e^{-ik}+2\mu^{*}&-2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ 2i\Delta_{1}e^{-i\varphi_{2}}\sin k&-t_{1}^{*}e^{-ik}-t_{2}^{*}e^{ik}-2\mu^{*}\end{array}\right)$ (B109)
$H(k)^{\dagger}$ $=\left(\begin{array}[]{cc}-t_{1}^{*}e^{ik}-t_{2}^{*}e^{-ik}-2\mu^{*}&2i\Delta_{1}e^{i\varphi_{1}}\sin k\\ -2i\Delta_{2}e^{-i\varphi_{2}}\sin k&t_{1}^{*}e^{-ik}+t_{2}^{*}e^{ik}+2\mu^{*}\end{array}\right)$ (B112)
@@
$\text{Tr}\big[H(k)^{2}\big]$ $=\text{Tr}\big[\{-H(k)\}^{2}\big]$ (B115)
$\text{Tr}\big[H(k)H(k)^{\dagger}\big]$ $=\text{Tr}\big[\{-H(k)\}\{-H^{\dagger}(k)\}\big]$ (B116)
@@
$\forall$ $\varphi_{1},\,\varphi_{2},\,\mu$ (B117)